[Brahms-l] [Fwd: Shift report 20031228 00:00-08:00]

From: Hironori Ito <hito@rcf.rhic.bnl.gov>
Date: Sun Dec 28 2003 - 10:14:03 EST

attached mail follows:


------------------------------------------------------------
Date: Sun Dec 28 07:33:52 2003, Shift: 00:00-08:00
Supervisor: Oana Ristea, with: Catalin Ristea
Log book#15 pages 17 - 17

Shift Summary:
------------------------------------------------------------
1. Starting the shift, the machine made some background studies for Phenix, so we had beam around 2 am. The beam was dumped at 5,30 am (yellow beam highly debunched). Waiting to refill, we got "RHIC CORRECTOR POWER SUPPLY/MAGNET PROBLEM - Making RHIC access to replace power supply bi8-th11".
2. Two data runs:
   9263 - D1 1/4A, D2 1/4A, D5 350A ~ 450k events
   9264 - D1 1/8A, D2 1/8A, D5 350A ~ 70k events
3. No critical new problems found for detectors.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Dec 28 10:13:59 2003

This archive was generated by hypermail 2.1.8 : Sun Dec 28 2003 - 10:14:19 EST